OpenStudy (anonymous):

How would I find the equation of an ellipse with center (4,3), vertex (4,-3) and a focus of (4,2)?

OpenStudy (anonymous):

I know the top part would be: (y-3)^2/... + (x-4)^2/... =1, but I don't know what is on the bottom.

OpenStudy (anonymous):

\[ c=1\\ b=6\\ b^2 = a^2 -c^2\\ 36 = a^2 -1\\ a^2 = 37\\ \frac {(x-4)^2}{37} +\frac {(y-3)^2}{36}=1 \]